OpenStudy (anonymous):

Algebra help? 5x + 2= 6x +11 Choose open; there is a variable true false Which value is a solution of the equation 5 - 4x= -3 Choose 0 2 -3 1/4 Which ordered pair is a solution of the equation y= x- 3 Choose -2,5 -5,2 2,5 5,2 Which ordered pair is a solution of the equation y=4x Choose 1,3 -1,-4 -4,-1 1,4 Which ordered pair is a solution of the equation y= -9x + 4 Choose 10,-86 -4, -58 6, -41 -6, 57 Which equation is an identity ? 11- (2v + 3) = -2v -8 5w + 8 - w= 6w - 2(w - 4) 7m - 2= 8m + 4 - m 8y + 9= 8y -3

pooja195 (pooja195):

1) Have you tried to solve the statement?

pooja195 (pooja195):

*expression

OpenStudy (anonymous):

Yes I have. I suck so bad at this stuff @pooja195

pooja195 (pooja195):

Well you don't suck you just need help and that's what this site is for! :) ok 1) Solve the equation see wether it is a true statment or not if it has a solution then it will be open

pooja195 (pooja195):

Solve for x and tell me what u get

OpenStudy (anonymous):

@pooja195 x=−9?

pooja195 (pooja195):

Yes, since there is one solution the statement is open

pooja195 (pooja195):

#2 solve for x

OpenStudy (anonymous):

@pooja195 thank god. x= 2 right?

pooja195 (pooja195):

Yes thats correct for #3 just plug in the given coordinates to see which makes a true statement Question: Which ordered pair is a solution of the equation y= x- 3 Choose -2,5 -5,2 2,5 5,2 Which of these makes a true statemnt 5= (-2)- 3 2= (-5)- 3 5= 2- 3 2= 5- 3

OpenStudy (anonymous):

@pooja195 Which equation is an identity ? I tried all of them, none of them are identity.. 11- (2v + 3) = -2v -8 5w + 8 - w= 6w - 2(w - 4) 7m - 2= 8m + 4 - m 8y + 9= 8y -3

pooja195 (pooja195):

which of the equations have the same solution is what it is basically asking so solve all